# Instruction Tuning for Large Language Models: A Survey

### v5

@@ 4.11 Others
(Conover et al., 2023a) is initialized with the pre-trained language model Pythia (12B) (Biderman et al., 2023), and fine-tuned on the instruction dataset databricks-dolly-15k888https://huggingface.co/datasets/databricks/databricks-dolly-15k, which contains 7 categories of NLP tasks such as text classification and information extraction. After fine-tuning, Dolly 2.0 (12B) outperforms Pythia (12B) on the EleutherAI LLM Evaluation Harness benchmark (Gao et al., 2021) by a large margin, and achieves comparable performances to GPT-NEOX (20B) (Black et al., 2022), which has two times more parameters compared to Dolly 2.0 (12B).
@@
(Almazrouei et al., 2023a) is a large language model trained by fine-tuning Falcon (40B) (Almazrouei et al., 2023b) on an English dialogue dataset, which contains 150 million tokens from the Baize dataset (Xu et al., 2023c), with an additional 5% of the data from the RefinedWeb dataset (Penedo et al., 2023). To reduce memory usage, the authors employed flash attention (Dao et al., 2022) and multi-query techniques. For evaluation, Falcon-Instruct (40B) achieved better performance on the Open LLM Leaderboard (Beeching et al., 2023)999https://huggingface.co/spaces/HuggingFaceH4
@@
is a large language model trained by fine-tuning the Starcoder Plus (15B) (Li et al., 2023f) on open-source instruction datasets including WizardLM (Xu et al., 2023a) and GPTeacher-General-Instruct101010https://github.com/teknium1/GPTeacher. For model inference, Minotaur supports a maximum context length of 18K tokens.
@@
is a large language model trained by fine-tuning LLaMA (13B) (Touvron et al., 2023a) on an instruction dataset, which contains over 300k instructions, sampled from GPTeacher111111https://github.com/teknium1/GPTeacher, CodeAlpaca (Chaudhary, 2023), GPT-4-LLM (Peng et al., 2023), Unnatural Instructions (Honovich et al., 2022), and BiologyPhysicsChemistry subsets in the Camel-AI (Li et al., 2023c). Responses are generated by GPT-4. For evaluations, Nous-Herme (13B) achieves comparable performances to GPT-3.5-turbo on multiple tasks like ARC challenge (Clark et al., 2018) and BoolQ (Clark et al., 2019).
@@
(Wang et al., 2023c) is a large language model trained by fine-tuning OPT (6.7B) (Zhang et al., 2022a) on a mixed instruction dataset, which contains FLAN V2 (Longpre et al., 2023), CoT (Wei et al., 2022), Dolly (Conover et al., 2023a), Open Assistant-1121212https://huggingface.co/datasets/OpenAssistant/oasst1, GPT4-Alpaca131313https://huggingface.co/datasets/vicgalle/alpaca-gpt4, Code-Alpaca (Chaudhary, 2023), and ShareGPT141414https://sharegpt.com/. After fine-tuning, TÜLU (6.7B) reaches on average 83% of ChatGPT’s performance and 68% of GPT-4’s performance.
@@
(YuLan-Chat-Team, 2023) is a language model trained by fine-tuning LLaMA (13B) (Touvron et al., 2023a) on a constructed bilingual dataset, which contains 250,000 Chinese-English instruction pairs. After fine-tuning, YuLan-Chat-13B achieves comparable results to the state-of-the-art open-source model ChatGLM (6B) (Du et al., 2022), and outperforms Vicuna (13B) (Chiang et al., 2023) on the English BBH3K (BBH3K is a subset of BBH benchmark (Srivastava et al., 2022)) dataset.
@@
151515https://txsun1997.github.io/blogs/moss.htmlis a bilingual dialogue language model, which aims to engage in multi-turn conversations and utilize various plugins, trained by fine-tuning on dialogue instructions. After fine-tuning, MOSS outperforms the backbone model and generates responses that better align with human preferences.
@@
161616https://github.com/jondurbin/airoborosis a large language model trained by fine-tuning LLAMA (13B) (Touvron et al., 2023a) on the Self-instruct dataset (Wang et al., 2022c). After fine-tuning, Airoboros significantly outperforms LLAMA (13B) (Touvron et al., 2023a) on all benchmarks and achieves highly comparable results to models fine-tuned specifically for certain benchmarks.
@@
5 Multi-modality Instruction Fine-tuning
@@ 5.1 Multi-modality Datasets
https://github.com/VT-NLP/MultiInstruct
@@
https://github.com/xiaoman-zhang/PMC-VQA
@@
https://github.com/OpenLAMM/LAMM
@@
https://vision-flan.github.io/
@@
https://github.com/FreedomIntelligence/ALLaVA
@@
https://sharegpt4v.github.io/
@@
(Zhang et al., 2023c) is a large-scale medical visual question-answering dataset that comprises 227k image-question pairs of 149k images, covering various modalities or diseases. The dataset can be used for both open-ended and multiple-choice tasks. The pipeline for generating the PMC-VQA dataset involves collecting image-caption pairs from the PMC-OA (Lin et al., 2023) dataset, using ChatGPT to generate question-answer pairs, and manually verifying a subset of the dataset for quality. The authors propose a generative-based model MedVInT for medical visual understanding by aligning visual information with a large language model. MedVInT pretrained on PMC-VQA achieves state-of-the-art performance and outperforms existing models on VQA-RAD (Lau et al., 2018) and SLAKE (Liu et al., 2021a) benchmarks, with 81.6% accuracy on VQA-RAD and 88.0% accuracy on SLAKE.
@@
5.2 Multi-modality Instruction Fine-tuning Models
@@
https://github.com/timothybrooks/instruct-pix2pix
@@
https://github.com/haotian-liu/LLaVA
@@
https://github.com/DAMO-NLP-SG/Video-LLaMA
@@
https://github.com/salesforce/LAVIS/tree/main/projects/instructblip
@@
https://github.com/Luodian/Otter
@@
https://github.com/open-mmlab/Multimodal-GPT
@@
(Zhang et al., 2023b) is a multimodal framework that enhances large language models with the ability to understand both visual and auditory content in videos. The architecture of Video-LLaMA consists of two branche encoders: the Vision-Language (VL) Branch and the Audio-Language (AL) Branch, and a language decoder (Vicuna (7B/13B) (Chiang et al., 2023), LLaMA (7B) (Touvron et al., 2023a), etc.). The VL Branch includes a frozen pre-trained image encoder (pre-trained vision component of BLIP-2 (Li et al., 2023d), which includes a ViT-G/14 and a pre-trained Q-former), a position embedding layer, a video Q-former and a linear layer. The AL Branch includes a pre-trained audio encoder (ImageBind (Girdhar et al., 2023)) and an Audio Q-former. Figure 8 shows the overall architecture of Video-LLaMA with Vision-Language Branch and Audio-Language Branch. The VL Branch is trained on the Webvid-2M (Bain et al., 2021) video caption dataset with a video-to-text generation task, and fine-tuned on the instruction-tuning data from MiniGPT-4 (Zhu et al., 2023b), LLaVA (Liu et al., 2023b) and VideoChat (Li et al., 2023e). The AL Branch is trained on video/image instru-caption data to connect the output of ImageBind to language decoder. After finetuning, Video-LLaMA can perceive and comprehend video content, demonstrating its ability to integrate auditory and visual information, understand static images, recognize common-knowledge concepts, and capture temporal dynamics in videos.
@@
6 Domain-specific Instruction Finetuning
@@
https://github.com/prakharguptaz/Instructdial
@@
https://github.com/BeyonderXX/InstructUIE
@@
https://github.com/amazon-science/instruction-tuning-for-absa
@@
https://github.com/facebookresearch/EditEval
@@
https://github.com/vipulraheja/coedit
@@
https://github.com/vishakhpk/creative-instructions
@@
https://huggingface.co/spaces/allen-eric/radiology-gpt
@@
https://github.com/Kent0n-Li/ChatDoctor
@@
https://github.com/SCIR-HI/Med-ChatGLM
@@
https://github.com/liutiedong/goat
@@
https://github.com/nlpxucan/WizardLM
@@ 6.3 Information Extraction
(Wang et al., 2023b) is a unified information extraction (IE) framework based on instruction tuning, which transforms IE tasks to the seq2seq format and solves them by fine-tuning 11B FlanT5 (Chung et al., 2022) on the constructed IT dataset. Figure 10 shows the overall architecture of InstructUIE. It introduces IE INSTRUCTIONS, a benchmark of 32 diverse information extraction datasets in a unified text-to-text format with expert-written instructions. Each task instance is delineated by four properties: task instruction, options, text, and output. Task instruction contains information such as the type of information to be extracted, the output structure format, and additional constraints or rules that need to be adhered to during the extraction process. Options refer to the output label constraints of a task. Text refers to the input sentence. Output is the sentence obtained by converting the original tags of the sample (e.g. "entity tag: entity span" for NER). In the supervised setting, InstructUIE performs comparably to BERT (Devlin et al., 2018) and outperforms the state-of-the-art and GPT3.5 (Brown et al., 2020a) in zero-shot settings.
@@ 6.6 Medical
(Li et al., 2023i) is based on the fine-tuned LLaMa-7B (Touvron et al., 2023a) model, utilizing the alpaca instruction dataset (Taori et al., 2023a) and the HealthCareMagic100k patient-doctor dialogue dataset. And prompt templates are designed for retrieving external knowledge databases, such as the Disease Database and Wikipedia retrieval, during doctor-patient conversations to obtain more accurate outputs from the model. The ChatDoctor significantly improves the model’s ability to comprehend patient needs and provide informed advice. By equipping the model with self-directed information retrieval from reliable online and offline sources, the accuracy of its responses is substantially improved.
@@
(Haochun Wang, 2023) is fine-tuned on the Chinese medical instruction dataset based on the ChatGLM-6B (Du et al., 2022) model. The instruction dataset comprises medically relevant question and answer pairs, created using the GPT 3.5 API and the Medical Knowledge Graph. This model improves the question-answering performance of ChatGLM (Du et al., 2022) in the medical field.
@@ 6.7 Arithmetic
(Liu and Low, 2023) is a fine-tuned LLaMA-7B (Touvron et al., 2023a) model based on instructions, which aims to solve arithmetic problems. It expresses arithmetic problems in the form of natural language question answering, such as "What is 8914/64?", by generating hundreds of instruction templates using ChatGPT (OpenAI, 2022). The model applies various techniques to enhance its adaptability to diverse question formats, such as randomly removing spaces between numbers and symbols in the arithmetic expression and replacing "*" with "x" or "times". The Goat model achieves state-of-the-art performance on the BIG-bench (Srivastava et al., 2022) arithmetic subtask. In particular, zero-shot Goat-7B matches or exceeds the accuracy achieved by the few-shot PaLM-540B (Chowdhery et al., 2022).
@@ 6.8 Code
(Luo et al., 2023) utilizes StarCoder 15B (Li et al., 2023f) as the foundation with complex instruction fine-tuning, by adapting the Evol-Instruct method (Xu et al., 2023a) to the domain of code. The training dataset is produced through iterative application of the Evol-Instruct technique on the Code Alpaca dataset (Taori et al., 2023b), which includes the following attributes for each sample: instruction, input, and expected output. For instance, when the instruction is "Amend the following SQL query to select distinct elements", the input is the SQL query, and the expected output is the generated answer. The WizardCoder outperforms all other open-source Code LLMs and even surpasses the largest closed LLMs, Anthropic’s Claude and Google’s Bard, on HumanEval and HumanEval+.

### v7

@@ 4.11 Others
(Conover et al., 2023a) is initialized with the pre-trained language model Pythia (12B) (Biderman et al., 2023), and fine-tuned on the instruction dataset databricks-dolly-15k999https://huggingface.co/datasets/databricks/databricks-dolly-15k, which contains 7 categories of NLP tasks such as text classification and information extraction. After fine-tuning, Dolly 2.0 (12B) outperforms Pythia (12B) on the EleutherAI LLM Evaluation Harness benchmark (Gao et al., 2021) by a large margin, and achieves comparable performances to GPT-NEOX (20B) (Black et al., 2022), which has two times more parameters compared to Dolly 2.0 (12B).
@@
(Almazrouei et al., 2023a) is a large language model trained by fine-tuning Falcon (40B) (Almazrouei et al., 2023b) on an English dialogue dataset, which contains 150 million tokens from the Baize dataset (Xu et al., 2023c), with an additional 5% of the data from the RefinedWeb dataset (Penedo et al., 2023). To reduce memory usage, the authors employed flash attention (Dao et al., 2022) and multi-query techniques. For evaluation, Falcon-Instruct (40B) achieved better performance on the Open LLM Leaderboard (Beeching et al., 2023)101010https://huggingface.co/spaces/HuggingFaceH4
@@
is a large language model trained by fine-tuning the Starcoder Plus (15B) (Li et al., 2023f) on open-source instruction datasets including WizardLM (Xu et al., 2023a) and GPTeacher-General-Instruct111111https://github.com/teknium1/GPTeacher. For model inference, Minotaur supports a maximum context length of 18K tokens.
@@
is a large language model trained by fine-tuning LLaMA (13B) (Touvron et al., 2023a) on an instruction dataset, which contains over 300k instructions, sampled from GPTeacher121212https://github.com/teknium1/GPTeacher, CodeAlpaca (Chaudhary, 2023), GPT-4-LLM (Peng et al., 2023), Unnatural Instructions (Honovich et al., 2022), and BiologyPhysicsChemistry subsets in the Camel-AI (Li et al., 2023c). Responses are generated by GPT-4. For evaluations, Nous-Herme (13B) achieves comparable performances to GPT-3.5-turbo on multiple tasks like ARC challenge (Clark et al., 2018) and BoolQ (Clark et al., 2019).
@@
(Wang et al., 2023d) is a large language model trained by fine-tuning OPT (6.7B) (Zhang et al., 2022a) on a mixed instruction dataset, which contains FLAN V2 (Longpre et al., 2023), CoT (Wei et al., 2022), Dolly (Conover et al., 2023a), Open Assistant-1131313https://huggingface.co/datasets/OpenAssistant/oasst1, GPT4-Alpaca141414https://huggingface.co/datasets/vicgalle/alpaca-gpt4, Code-Alpaca (Chaudhary, 2023), and ShareGPT151515https://sharegpt.com/. After fine-tuning, TÜLU (6.7B) reaches on average 83% of ChatGPT’s performance and 68% of GPT-4’s performance.
@@
(YuLan-Chat-Team, 2023) is a language model trained by fine-tuning LLaMA (13B) (Touvron et al., 2023a) on a constructed bilingual dataset, which contains 250,000 Chinese-English instruction pairs. After fine-tuning, YuLan-Chat-13B achieves comparable results to the state-of-the-art open-source model ChatGLM (6B) (Du et al., 2022), and outperforms Vicuna (13B) (Chiang et al., 2023) on the English BBH3K (BBH3K is a subset of BBH benchmark (Srivastava et al., 2022a)) dataset.
@@
161616https://txsun1997.github.io/blogs/moss.htmlis a bilingual dialogue language model, which aims to engage in multi-turn conversations and utilize various plugins, trained by fine-tuning on dialogue instructions. After fine-tuning, MOSS outperforms the backbone model and generates responses that better align with human preferences.
@@
171717https://github.com/jondurbin/airoborosis a large language model trained by fine-tuning LLAMA (13B) (Touvron et al., 2023a) on the Self-instruct dataset (Wang et al., 2022c). After fine-tuning, Airoboros significantly outperforms LLAMA (13B) (Touvron et al., 2023a) on all benchmarks and achieves highly comparable results to models fine-tuned specifically for certain benchmarks.
@@
5 Multi-modality Instruction Tuning
@@ 5.1 Multi-modality Datasets
1 https://github.com/VT-NLP/MultiInstruct
@@
2 https://github.com/xiaoman-zhang/PMC-VQA
@@
3 https://github.com/OpenLAMM/LAMM
@@
4 https://vision-flan.github.io/
@@
5 https://github.com/FreedomIntelligence/ALLaVA
@@
6 https://sharegpt4v.github.io/
@@
(Zhang et al., 2023c) is a large-scale medical visual question-answering dataset that comprises 227k image-question pairs of 149k images, covering various modalities or diseases. The dataset can be used for both open-ended and multiple-choice tasks. The pipeline for generating the PMC-VQA dataset involves collecting image-caption pairs from the PMC-OA (Lin et al., 2023b) dataset, using ChatGPT to generate question-answer pairs, and manually verifying a subset of the dataset for quality. The authors propose a generative-based model MedVInT for medical visual understanding by aligning visual information with a large language model. MedVInT pretrained on PMC-VQA achieves state-of-the-art performance and outperforms existing models on VQA-RAD (Lau et al., 2018) and SLAKE (Liu et al., 2021a) benchmarks, with 81.6% accuracy on VQA-RAD and 88.0% accuracy on SLAKE.
@@
5.2 Multi-modality Instruction Tuning Models
@@
1 https://github.com/timothybrooks/instruct-pix2pix
@@
2 https://github.com/haotian-liu/LLaVA
@@
3 https://github.com/DAMO-NLP-SG/Video-LLaMA
@@
4 https://github.com/salesforce/LAVIS/tree/main/projects/instructblip
@@
5 https://github.com/Luodian/Otter
@@
6 https://github.com/open-mmlab/Multimodal-GPT
@@
(Zhang et al., 2023b) is a multimodal framework that enhances large language models with the ability to understand both visual and auditory content in videos. The architecture of Video-LLaMA consists of two branche encoders: the Vision-Language (VL) Branch and the Audio-Language (AL) Branch, and a language decoder (Vicuna (7B/13B) (Chiang et al., 2023), LLaMA (7B) (Touvron et al., 2023a), etc.). The VL Branch includes a frozen pre-trained image encoder (pre-trained vision component of BLIP-2 (Li et al., 2023d), which includes a ViT-G/14 and a pre-trained Q-former), a position embedding layer, a video Q-former and a linear layer. The AL Branch includes a pre-trained audio encoder (ImageBind (Girdhar et al., 2023)) and an Audio Q-former. Figure 8 shows the overall architecture of Video-LLaMA with Vision-Language Branch and Audio-Language Branch. The VL Branch is trained on the Webvid-2M (Bain et al., 2021) video caption dataset with a video-to-text generation task, and fine-tuned on the instruction tuning data from MiniGPT-4 (Zhu et al., 2023b), LLaVA (Liu et al., 2023b) and VideoChat (Li et al., 2023e). The AL Branch is trained on video/image instru-caption data to connect the output of ImageBind to language decoder. After finetuning, Video-LLaMA can perceive and comprehend video content, demonstrating its ability to integrate auditory and visual information, understand static images, recognize common-knowledge concepts, and capture temporal dynamics in videos.
@@
6 Domain-specific Instruction Tuning
@@
1 https://github.com/prakharguptaz/Instructdial
@@
2 https://github.com/BeyonderXX/InstructUIE
@@
3 https://github.com/amazon-science/instruction-tuning-for-absa
@@
4 https://github.com/facebookresearch/EditEval
@@
5 https://github.com/vipulraheja/coedit
@@
6 https://github.com/vishakhpk/creative-instructions
@@
7 https://huggingface.co/spaces/allen-eric/radiology-gpt
@@
8 https://github.com/Kent0n-Li/ChatDoctor
@@
9 https://github.com/SCIR-HI/Med-ChatGLM
@@
10 https://github.com/liutiedong/goat
@@
11 https://github.com/nlpxucan/WizardLM
@@ 6.3 Information Extraction
(Wang et al., 2023c) is a unified information extraction (IE) framework based on instruction tuning, which transforms IE tasks to the seq2seq format and solves them by fine-tuning 11B FlanT5 (Chung et al., 2022) on the constructed IT dataset. Figure 10 shows the overall architecture of InstructUIE. It introduces IE INSTRUCTIONS, a benchmark of 32 diverse information extraction datasets in a unified text-to-text format with expert-written instructions. Each task instance is delineated by four properties: task instruction, options, text, and output. Task instruction contains information such as the type of information to be extracted, the output structure format, and additional constraints or rules that need to be adhered to during the extraction process. Options refer to the output label constraints of a task. Text refers to the input sentence. Output is the sentence obtained by converting the original tags of the sample (e.g. "entity tag: entity span" for NER). In the supervised setting, InstructUIE performs comparably to BERT (Devlin et al., 2018) and outperforms the state-of-the-art and GPT3.5 (Brown et al., 2020a) in zero-shot settings.
@@ 6.6 Medical
(Li et al., 2023j) is based on the fine-tuned LLaMa-7B (Touvron et al., 2023a) model, utilizing the alpaca instruction dataset (Taori et al., 2023a) and the HealthCareMagic100k patient-doctor dialogue dataset. And prompt templates are designed for retrieving external knowledge databases, such as the Disease Database and Wikipedia retrieval, during doctor-patient conversations to obtain more accurate outputs from the model. The ChatDoctor significantly improves the model’s ability to comprehend patient needs and provide informed advice. By equipping the model with self-directed information retrieval from reliable online and offline sources, the accuracy of its responses is substantially improved.
@@
(Wang et al., 2023a) is fine-tuned on the Chinese medical instruction dataset based on the ChatGLM-6B (Du et al., 2022) model. The instruction dataset comprises medically relevant question and answer pairs, created using the GPT 3.5 API and the Medical Knowledge Graph. This model improves the question-answering performance of ChatGLM (Du et al., 2022) in the medical field.
@@ 6.7 Arithmetic
(Liu and Low, 2023) is a fine-tuned LLaMA-7B (Touvron et al., 2023a) model based on instructions, which aims to solve arithmetic problems. It expresses arithmetic problems in the form of natural language question answering, such as "What is 8914/64?", by generating hundreds of instruction templates using ChatGPT (OpenAI, 2022). The model applies various techniques to enhance its adaptability to diverse question formats, such as randomly removing spaces between numbers and symbols in the arithmetic expression and replacing "*" with "x" or "times". The Goat model achieves state-of-the-art performance on the BIG-bench (Srivastava et al., 2022a) arithmetic subtask. In particular, zero-shot Goat-7B matches or exceeds the accuracy achieved by the few-shot PaLM-540B (Chowdhery et al., 2022).
@@ 6.8 Code
(Luo et al., 2023) utilizes StarCoder 15B (Li et al., 2023f) as the foundation with complex instruction tuning, by adapting the Evol-Instruct method (Xu et al., 2023a) to the domain of code. The training dataset is produced through iterative application of the Evol-Instruct technique on the Code Alpaca dataset (Taori et al., 2023b), which includes the following attributes for each sample: instruction, input, and expected output. For instance, when the instruction is "Amend the following SQL query to select distinct elements", the input is the SQL query, and the expected output is the generated answer. The WizardCoder outperforms all other open-source Code LLMs and even surpasses the largest closed LLMs, Anthropic’s Claude and Google’s Bard, on HumanEval and HumanEval+.
@@ 8 Evaluation, Analysis and Criticism
8.1 Close-ended Evaluations
It is widely accepted among researchers that general-purpose models must demonstrate proficiency in certain core tasks before they can effectively generalize to meet diverse real-world needs. Close-ended evaluations help achieve this objective, often involving multiple-choice questions to assess the performance of LLMs. Below are 6 widely used close-ended evaluations:
(1) MMLU.
Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2020a) consists of 14079 questions covering 57 tasks including elementary mathematics, US history, computer science, law, and more. The wide range of subjects and complex questions make MMLU suitable for testing the model’s language comprehension and decision-making capabilities.
(2) MATH and (3) GSM8K.
MATH Hendrycks et al. (2021) and GSM8K Cobbe et al. (2021) are two distinct mathematical datasets utilized for evaluating different aspects of model capabilities. The MATH Hendrycks et al. (2021) dataset comprises 12,500 complex competition-level mathematics problems, primarily designed to assess the ability of models to tackle challenging and advanced mathematical questions typically encountered at the college level. Conversely, the GSM8K Cobbe et al. (2021) dataset contains 8,500 high-quality elementary school math problems, aimed at testing the basic mathematical reasoning abilities of models.
(4) BBH.
BBH, short for BIG-Bench Hard Suzgun et al. (2022a), is a subset of the BIG-Bench Srivastava et al. (2022b) dataset comprising 23 challenging tasks. These tasks were selected because they consistently proved too difficult for current large language models to handle effectively. Requiring complex, multi-step reasoning, the BBH dataset is primarily utilized to assess the general reasoning capabilities of models, testing their ability to navigate and solve intricate problems.
(5) HumanEval (Coding).
HumanEval Chen et al. (2021a) consists of 164 programming problems, including language comprehension, algorithms, and simple mathematics, with some comparable to simple software interview questions. The primary purpose of this dataset is to assess the ability of models to generate correct programs based on provided docstrings.
(6) IFEval.
IFEval Zhou et al. (2023b) consists of 500 prompts, each containing specific instructions like "write an article with more than 800 words" or "enclose your response in double quotation marks." This dataset is used to test the ability of large language models to accurately follow given instructions.
8.2 GPT-Based Evaluations
GPT-based evaluations refer to methods that leverage GPT-4 OpenAI (2023) to assess the output of LLMs. There are three primary reasons for this approach: (1) Efficiency – Manually reviewing numerous LLM outputs can be labor-intensive, whereas GPT-4 can evaluate large-scale responses quickly, saving both time and effort; (2) Reliable Benchmark – As one of the most advanced models available, GPT-4 provides a dependable benchmark, allowing researchers to compare the performance of different LLMs against a high standard; and (3) Enhanced Capability – With improved comprehension and reasoning over previous models, GPT-4 is better suited to analyze subtle aspects of language generation and handle complex outputs from other LLMs. In the following, we detail 4 commonly accepted GPT-based evaluations:
(1) AlpacaEval.
AlpacaEval Li et al. (2023h) is an automated evaluation metric leveraging LLMs, consisting of 805 instructions selected to reflect typical user interactions from the Alpaca web demo181818https://crfm.stanford.edu/2023/03/13/alpaca.html. Specifically, for each instruction, both a baseline model (currently GPT-4 turbo OpenAI (2023)) and the model under evaluation generate responses. A GPT-4 turbo-based evaluator then conducts a head-to-head comparison of these responses, determining the probability of favoring the evaluated model. The win rate is calculated as the expected probability that the evaluator prefers the evaluated model’s response across the 805 instructions, serving as a key metric for assessing the performance of the evaluated LM chatbot.
(2) Length-Controlled AlpacaEval.
Length-Controlled AlpacaEval Dubois et al. (2024) is a variation of the AlpacaEval Li et al. (2023h) evaluation metric, designed to minimize length bias, as the original AlpacaEval tends to favor models that produce longer responses. To achieve this goal, Dubois et al. (2024) first fit a generalized linear model to predict the annotator’s (GPT-4’s) preference based on three factors: (1) the instruction, (2) the model used, and (3) the length difference between the baseline and the model’s output. Then, by conditioning the length difference to 0, Dubois et al. (2024) can obtain the length-controlled preference. This idea, which predicts the outcome while conditioning on the length difference (mediator), is a common technique in statistical inference, and by introducing it, Length-Controlled AlpacaEval increases the Spearman correlation with LMSYS’ Chatbot Arena from 0.94 to 0.98.
(3) MT-Bench.
Currently, close-ended evaluations only measure LLMs’ core capability on a confined set of tasks, such as MMLU Hendrycks et al. (2020a) for multi-choice decisions, without adequately assessing its alignment with human preference in open-ended tasks, such as the ability to adhere to instructions in multi-turn dialogues accurately. To alleviate this issue, Zheng et al. (2023) introduced MT-Bench, which comprises 80 high-quality multi-turn questions designed to assess LLMs’ capability in multi-turn conversations and instruction-following, with evaluations conducted using GPT-4. MT-Bench is meticulously crafted to cover eight common tasks: writing, roleplay, extraction, reasoning, math, coding, knowledge I (STEM), and knowledge II (humanities/social sciences). For alignment, GPT-4 achieves over 80% agreement, comparable to the level of agreement among humans, making it a more reliable choice for a public benchmark.
(4) WildBench.
Although the above evaluations are effective, they have notable limitations in task composition and skill coverage. For example, MT-Bench Hendrycks et al. (2020a) includes only 80 test instructions, while AlpacaEval Li et al. (2023h) features many straightforward tasks, such as “What is the capital of Australia?” To address this issue, Lin et al. (2024) introduced WildBench, comprising 1,024 test instructions carefully curated from extensive human-chatbot conversation logs. WildBench draws directly from real-world user interactions, featuring numerous challenging tasks, such as coding and math problem-solving. These tasks frequently demand critical thinking, making WildBench significantly more difficult than other benchmarks. WildBench utilizes two metrics: WB-Reward for pairwise comparisons and WB-Score for individual assessments. Both metrics show strong alignment with human evaluations, with Pearson correlations of 0.98 for WB-Reward and 0.95 for WB-Score when compared to the human-voted ratings.